OpenStudy (anonymous):

Given f(x) = 2x3 + 5, what is f '(2)?

OpenStudy (anonymous):

is it 2*3+5?

OpenStudy (anonymous):

2x^3 two x squared

OpenStudy (anonymous):

if it is 2x cube(3) +5 f(2)=2(2)3+5 =8+5=13

OpenStudy (anonymous):

yeah i meant two x cubed. but that'a not the derivative "/

OpenStudy (anonymous):

oh in the cube its wrong.16+5=21

OpenStudy (anonymous):

is that a derivative?

OpenStudy (anonymous):

yes! :)

OpenStudy (anonymous):

=d2x3+d5 =3*2x to the power of 3-1+0 =6x2

OpenStudy (anonymous):

24! thanksss